OpenStudy (cheesecakekitten):

WILL MEDAL AND FAN!!! A rectangle is (2x + 5) long and (x^2 - 3x + 1) wide. What is its area?

OpenStudy (cheesecakekitten):

@madiimay4music

OpenStudy (cheesecakekitten):

@Mehek14 @ParthKohli

OpenStudy (cheesecakekitten):

@imqwerty @Jaynator495

OpenStudy (cheesecakekitten):

@NetflixAndNo

OpenStudy (cheesecakekitten):

@AloneS

OpenStudy (cheesecakekitten):

SOMEONE HELP

NetflixAndNo (netflixandno):

I can't figure it out... I'm sorry :(

OpenStudy (cheesecakekitten):

Thanks for trying

OpenStudy (tygrr321):

@YoungStudier can help u

OpenStudy (youngstudier):

@Awolflover1

OpenStudy (tygrr321):

nvm lel

OpenStudy (cheesecakekitten):

;-;

OpenStudy (anonymous):

In rectangle, the distance around the outside of the rectangle is known as perimeter. A rectangle is 2-dimensional; however, perimeter is 1-dimensional and is measured in linear units such as feet or meter etc. The perimeter of a rectangle is the total length of all the four sides. Perimeter of rectangle = 2L + 2W. Example 1: Rectangle has the length 13 cm and width 8 cm. solve for perimeter of rectangle. Solution: Given that: Length (l) = 13 cm Width (w) = 8 cm Perimeter of the rectangle = 2(l + w) units P = 2(13 + 8) P = 2 (21) P = 42 Thus, the perimeter of the rectangle is 42 cm. Example 2: If a rectangle's length is 2x + 1 and its width is 2x – 1. If its area is 15 cm2, what are the rectangle's dimensions and what is its perimeter? Solution: We know that the dimensions of the rectangle in terms of x: l = 2x + 1 w = 2x – 1

OpenStudy (anonymous):

Since the area of a rectangle is given by: A = l * w We can substitute the expressions for length and width into the equation for area in order to determine the value of x. A = l * w 15 = (2x + 1) (2x -1) 15 = 4x2 – 1 16 = 4x2 x = ±2 Note that the value of x must be positive and therefore in our case, the value of x is 2. And now we have: l = 5 cm w = 3 cm Therefore, the dimensions are 5cm and 3cm. Now, substituting these values in the formula for perimeter, we will get P = 2l + 2w P = 2(5)+2(3) P = 10+6 P = 16 cm Example 3: Find the area and the perimeter of a rectangle whose length is 24 m and width is 12m? Solution: Given that: length = L = 24m width = W = 12m Area of a rectangle: A = L × W A = 24 × 12 A = 188 m2 Perimeter of a rectangle: P = 2L + 2W P = 2(24) + 2(12) P = 48 + 24 P = 72 m Example 4: Find the area and perimeter of a rectangle whose breadth is 4 cm and the height 3 cm. Solution: Area = b×h = 4×3 = 12 cm2. Perimeter = 2(b) + 2(h) = 2(4) + 2(3) = 8 + 6 = 14. Example 5: Calculate the perimeter of the rectangle whose length is 18cm and breadth 7cm Solution: Given that: L = 18 cm B = 7 cm Perimeter of rectangle = 2(length + breadth) P = 2 (L + B) P = 2 (18 + 7) P = 50 cm Example 6: Find the perimeter of rectangle whose length is 6 inches and width is 4 inches. Solution: P = 2(L + B) P = 2(6 + 4) P = 20 in Example 7: A boy walks 5 times around a park. If the size of the park is 100m by 50m, find the distance the boy has walked. If he walks 100m in 5 minutes, how long will it take for him in total? Solution: Given that: Length = L = 100m Width = W = 50m Rounds = 5 Time per 100m = 5minutes. Perimeter of the park: P = 2 L + 2 W. P = 2 × 100 + 2 × 50 P = 200 + 100 P = 300 m Total distance walked = 5 × Perimeter of the park. = 5 × 300 = 1500 meters Total time taken = Total distance walked × time taken to walk 1m. = 1500 × 5/100 = 75minutes or 1hr 15minutes Online Perimeter Calculator Web-Formulas.com © 2016 | Contact us | Terms of Use | Privacy Policy | Yellow Sparks Network Web Formulas

OpenStudy (madiimay4music):

@CheesecakeKitten

OpenStudy (madiimay4music):

your answer is a